# BlueME: Robust Underwater Robot-to-Robot Communication Using Compact Magnetoelectric Antennas

### v2

@@ Abstract
We present the design, development, and experimental validation of BlueME, a compact magnetoelectric (ME) antenna array system for underwater robot-to-robot communication. BlueME employs ME antennas operating at their natural mechanical resonance frequency to efficiently transmit and receive very-low-frequency (VLF) electromagnetic signals underwater. We outline the design, simulation, fabrication, and integration of the proposed system on low-power embedded platforms focusing on portable and scalable applications. For performance evaluation, we deployed BlueME on an autonomous surface vehicle (ASV) and a remotely operated vehicle (ROV) in open-water field trials. Our tests demonstrate that BlueME maintains reliable signal transmission at distances beyond 200 meters while consuming only 1 watt of power. Field trials show that the system operates effectively in challenging underwater conditions such as turbidity, obstacles, and multipath interference– that generally affect acoustics and optics. Our analysis also examines the impact of complete submersion on system performance and identifies key deployment considerations. This work represents the first practical underwater deployment of ME antennas outside the laboratory, and implements the largest VLF ME array system to date. BlueME demonstrates significant potential for marine robotics and automation in multi-robot cooperative systems and remote sensor networks.
@@ 1 Introduction
Traditional electromagnetic and radio frequency (RF) signals are severely attenuated underwater [1, 10], limiting their effective communication range to only a few meters. While propagation is acceptable with very low frequencies (VLFs) [11], it typically comes at the cost of high transmission power requirements and impractically large antenna sizes [12, 13]. In seawater, electromagnetic signal absorption ranges from 1 to 10 dB/meter depending on conductivity [14]. This high attenuation severely limits the viability of RF communications in practical applications [15].
Acoustic communication has been the most commonly used modality for underwater robot communications, offering omnidirectional transmission and acceptable signal attenuation over longer distances [16, 17]. Low-frequency long-range systems are used in ocean basin communication [18], while high-frequency short-range systems are preferred for robot positioning and data transfer [19, 20]. Acoustic communications suffer from Doppler effects, phase and magnitude fluctuations, and multi-path interference [21, 22, 23], compromising reliability and efficiency. Acoustic signals can be disruptive to marine ecosystems as well, limiting their long-term deployments. Optical communication systems (e.g., visible light, blue-green lasers) offer higher data rates [24, 25, 26], but are limited by line-of-sight requirements. They are susceptible to biofouling [27], a process where microorganisms and biological buildup degrade optical sensors and signal quality over time [28].
Magnetoelectric (ME) antennas offer a promising solution to overcome the limitations of acoustic and optical communication in underwater environments [29, 11, 30, 31, 32]. By leveraging the magnetoelectric effect, these antennas efficiently transmit and receive very-low-frequency (VLF) electromagnetic signals with a more compact design than traditional electrically small antennas (ESAs). This efficiency, combined with their resilience to multi-path interference, non-line-of-sight requirements, and Doppler effects, makes ME antennas well-suited for space- and power-constrained underwater applications.
In this paper, we present the design, development, and validation of BlueME, a novel magnetoelectric (ME) antenna system for real-time data transmission between underwater robots. As shown in Fig. 1, we use a compact array of ME antennas integrated into pressure-compensated enclosures for real-time data transmission and retrieval between robots and sensors deployed underwater. While previous ME-based underwater communication efforts have primarily focused on basic connectivity tests in controlled water-tank environments [29, 33, 34, 11], we validate BlueME in a natural lake, demonstrating its potential as a deployable communication system for marine robotics applications.
@@
We present a seamless integration of BlueME into pressure-compensated enclosures to ensure reliable underwater operations. The enclosed system is compact and portable for use on low-power embedded devices.
We validate the system through simulation as well as open-water experiments, demonstrating its effectiveness for underwater communication. The system performs reliably when fully submerged and is unaffected by turbidity, line-of-sight obstacles, and shallow-water interference, which limit acoustic and optical systems.
Our field deployments reveal that we can achieve high-fidelity communications between mobile robots for up to meters with a power footprint of only watt. These capabilities demonstrate its potential for robust and low-latency communication underwater.
@@ 2.1 Underwater Communication Technologies
Optical communication offers higher data rates and lower latency compared to acoustic methods [24, 25]. Notable systems include the BlueComm system, achieving 20 Mbps over distances up to 200 m [42]; the Ambalux system, providing 10 Mbps over 40 m [25]; and LUMA modems, supporting 10 Mbps across distances up to 50 m [43]. Underwater visible light communication (UVLC) has also been explored for extended ranges using wavelengths between 450–550 nm, which experience lower attenuation in water [26, 28]. However, optics are limited by line-of-sight requirements and are susceptible to scattering, absorption due to turbidity, and biofouling effects that degrade performance over time [44, 45, 46].
Radio frequency (RF) electromagnetic communication underwater has historically been challenging due to high attenuation of EM waves in conductive media like seawater [14, 15]. Attenuation increases with frequency, limiting effective communication to short distances [15]. However, RF communication offers advantages such as the ability to cross air-water and water-sediment boundaries, operation in turbid conditions, and non-line-of-sight capabilities [10, 47]. Recent research has revisited underwater RF communication using very low frequency (VLF, 3–30 kHz) and ultra-low frequency (ULF, 300–3000 Hz) bands to mitigate attenuation [29, 10]. Despite these advances, practical deployment is hindered by the need for large antennas and/or high transmission power, making it less suitable for mobile underwater applications [48, 12].
@@ 2.2 ME Antennas for Low-Power Communication
However, challenges remain in designing and fabricating compact, low-power ME antennas suitable for integration with mobile robotic platforms [54, 55]. Key issues include optimizing antenna performance, developing feasible deployment strategies, and characterizing communication properties in dynamic underwater conditions. This work aims to advance the field by developing a novel ME antenna array system tailored for marine robotics applications, addressing current gaps in practical deployment and performance evaluation.
@@ 3.1 ME Antenna Design: Theory and Simulation
To validate this theoretical analysis, we use COMSOL Multiphysics [61] to simulate a three-layer ME structure. As shown in Fig. 3, we evaluate the displacement magnitude of our antenna with a Metglas-PZT-Metglas thickness of -- m. By simulation, we predict the eigenfrequency for the fundamental displacement mode to be kHz.
Antenna Enclosure and Array Assembly. As shown in Fig. 4, each ME antenna is housed in a custom 3D-printed enclosure. The upper shell of the enclosure contains a pogo pin and two vertical AWG copper wires; the pogo pin is bonded to a designated hole on the antenna using a UV-curable cyanoacrylate adhesive. The AWG wires from the antenna are twisted and soldered to the AWG wires, which extend through the lower shell. The two shells are then enclosed and arranged in a grid within a frame that includes / inch N52 neodymium magnets oriented north-to-south to maintain the required magnetic bias. The receiver and transmitter antennas are connected in series and parallel configurations, respectively.
3.2 System Integration
@@ 4 System Integration & Experimental Setup
Our experimental analyses include (i) evaluation of the BlueME system’s magnetoelectric (ME) antenna communication characteristics, and (ii) performance assessment through robotics field trials. First, we validate the proposed antenna’s functionality using empirical data and simulation results, providing a basis for further investigation into optimal modulation techniques. We then integrate two BlueME systems on an autonomous surface vehicle (ASV) and an underwater remotely operated vehicle (ROV) to evaluate communication range and robustness, as illustrated in Fig. 6. To this end, we conduct a series of freshwater trials to demonstrate the feasibility of the proposed BlueME system for robot-to-robot communication.
@@ 4.1 Setup: Measuring Antenna Impedance
An impedance sweep from to was performed using a DAD3 board connected to an impedance measurement attachment and a BNC adapter. In these tests, the antennas were connected via short wires to a terminal block and BNC oscilloscope probes. The impedance magnitude and phase were recorded for antennas individually.
After individual tests, the antennas were wired together to form the transmitter and receiver arrays. The transmitter antennas were connected in parallel, while the receiver antennas were connected in series. The arrays were then tested as a whole, with impedance measurements taken under the same conditions as the individual antennas. The goal was to fabricate the ME antennas with sufficiently tight tolerances to validate that wiring the antennas in parallel/series effectively increases bandwidth and/or resonant response [62].
@@ 4.2 Setup: Underwater Communication Range
We first conducted proof-of-concept experiments by evaluating the impedance of the proposed system in a water tank measuring m (L) m (W) m (D). The range tests were conducted using two ME antenna arrays in a freshwater lake. The receiver array was mounted on an underwater ROV (BlueROV2) and positioned at a depth of approximately below the surface, whereas the transmitter array was mounted on an ASV (BlueBoat) with the transmitter submerged approximately below the surface. The transmitter array was driven by a Class-D amplifier powered by a supply, which was generated using a boost converter connected to a LiFePO4 battery. The input signal to the amplifier varied from to . A small monitoring board was placed in line with the amplifier output, consisting of current sense resistors and an AC-coupled : voltage divider to measure the current through and voltage across the antenna array.
@@
For the range tests, the transmitter array conducted frequency sweeps from to at varying drive strengths. The receiver array collected data continuously, using the FFT peak hold feature of the WaveForms software to record the received signal amplitude across the frequency range. Tests were performed at distances of , , , and , with the boat’s position verified via GPS with an accuracy of approximately .
5 Performance Evaluation
@@ 5.1 Impedance Measurement Results
5.2 Underwater Communication Performance
For the range tests, we initially anticipated that near-field behavior would dominate the antenna response, causing rapid signal attenuation with distance. However, the results demonstrated effective communication at distances from up to , indicating that far-field effects and alternative propagation paths played a significant role underwater. Note that our instruments and open-water logistics limited the experiments within this range; thus, BlueME can potentially communicate beyond .
This performance can be attributed to several factors. First, the electromagnetic wavelength in water (see 6.1) at is approximately , much shorter than in air due to water’s higher permittivity and conductivity. The wavelength in a conductive medium is given by , where is the phase constant. In conductive media, is influenced by permittivity, permeability, conductivity, and angular frequency. The reduced wavelength improves the radiation efficiency of electrically small antennas [11]. Additionally, the ME antennas, when configured in arrays, act collectively within each other’s near fields, effectively forming a larger magnetic dipole with increased radiation [11].
Moreover, we found that the rate of attenuation decreased at greater distances. This may be due to electromagnetic waves propagating along paths with lower attenuation (e.g., the lake bed, surface wave) [63] or due to reflections at shallow water surfaces [15]. Unlike acoustic propagation, electromagnetic waves’ significantly higher propagation velocity in water results in wavelengths of orders of magnitude longer than acoustic waves at comparable frequencies. Consequently, typical path length differences produce smaller phase differences between multipath components, rendering interference effects negligible. Additionally, reflections from obstacles and natural features are minimal since objects of size comparable to or smaller than the electromagnetic wavelength produce weak scattering, with incident waves primarily diffracting around or penetrating through such objects rather than producing strong reflections.
6 Analysis & Discussion
@@ 6.1 Analyzing the Theory of Operation
The performance characteristics of ME antenna arrays submerged underwater are not extensively explored in the literature. As such, we make the following set of observations in our experimental analyses.
Wavelength Changes. The antenna’s observed signal drop-off appears to change across distances. This may be explained by the change in electromagnetic wavelength from air to fully submerged in water, which is also a more conductive medium. To this end, we calculate the expected wavelength of an electromagnetic plane wave in an underwater environment. As shown by [14], the electromagnetic propagation constant can be calculated as:
@@
Although the relative permittivity of water is significantly higher than air (:), the substantial decrease in wavelength (from in air as opposed to in water) results in a net increase in radiation resistance per antenna by a factor of approximately . This leads to an increase in the total radiated power per antenna by a similar factor.
@@
Increasing the number of antennas on the receiver side also increases the receiver sensitivity linearly, assuming all antennas are identical, i.e., a factor of improvement in the overall link budget. Consequently, using an array of antennas underwater compared to a single antenna in air theoretically yields an overall link budget improvement by a factor of , which corresponds to an increase of approximately dB – a factor not specifically discussed in the literature [11, 48, 29, 33, 34]. In practical scenarios, however, resonant frequency mismatches and increased mutual coupling at closer antenna spacings and higher drive strengths will degrade phase coherence within the transmitter array [62], reducing the actual achievable improvement from the array configuration.

### v4

@@ Abstract
We present the design, development, and experimental validation of BlueME, a compact magnetoelectric (ME) antenna array system for underwater robot-to-robot communication. BlueME employs ME antennas operating at their natural mechanical resonance frequency to efficiently transmit and receive very-low and low-frequency (VLF/LF) electromagnetic signals underwater. We detail the system’s design, simulation, fabrication, and integration onto low-power embedded platforms, emphasizing portability and scalability. For performance evaluation, we deployed BlueME on mobile robot platforms in freshwater (lake) and saltwater (ocean) trials. Our tests demonstrate reliable signal transmission and detection using the BlueME antenna system at distances beyond 700 meters while consuming less than 10 watts of power. Ocean trials demonstrate that the system operates effectively under challenging conditions, such as turbidity, obstacles, and multipath interference, which are factors that typically degrade acoustic and optical methods. Our analysis also examines the impact of complete submersion on system performance and identifies key deployment considerations. This work represents the first practical underwater deployment of ME antennas outside the laboratory and implements the largest VLF/LF ME array system to date. BlueME demonstrates significant potential for marine robotics and automation in multi-robot cooperative systems and remote sensor networks.
@@ 1 Introduction
Conventional electromagnetic and radio frequency (RF) signals are severely attenuated underwater [1, 10], limiting their effective communication range to only a few meters. While propagation improves at very-low and low-frequencies (VLFs/LFs) [11], it typically comes at the cost of high transmission power and impractically large antenna sizes [12, 13]. In seawater, electromagnetic signal attenuation ranges from approximately 1 to 10 dB per meter depending on both frequency and conductivity [14]. This high attenuation severely limits the viability of RF communication for marine robotics applications [15].
Acoustic communication has been the most commonly used modality for underwater robot communications, offering omnidirectional transmission and acceptable signal attenuation over longer distances [16, 17]. Long-range low-frequency systems are used in ocean-based communication [18], while short-range high-frequency systems are preferred for robot positioning and data transfer [19, 20]. Acoustic communications suffer from Doppler effects, phase and magnitude fluctuations, and multi-path interference [21, 22, 23], compromising reliability and efficiency. Acoustic signals can be disruptive to marine ecosystems as well, limiting their long-term deployments. Optical communication systems (e.g., visible light, blue-green lasers) offer higher data rates [24, 25, 26], but are limited by line-of-sight requirements. They are susceptible to biofouling [27], a process where microorganisms and biological buildup degrade optical sensors and signal quality over time [28].
Magnetoelectric (ME) antennas offer a promising solution to overcome the limitations of acoustic and optical communication in underwater environments [29, 11, 30, 31, 32]. By leveraging the magnetoelectric effect, these antennas efficiently transmit and receive electromagnetic signals with a more compact design than traditional electrically small antennas (ESAs) [33]. This efficiency, combined with their resilience to multi-path interference, non-line-of-sight requirements, and Doppler effects, makes ME antennas well-suited for space- and power-constrained underwater applications.
In this paper, we present the design, development, and validation of BlueME, a novel magnetoelectric (ME) antenna system for real-time data transmission between underwater robots. As shown in Fig. 1, we use a compact array of ME antennas integrated into pressure-compensated enclosures for real-time data transmission and retrieval between robots and sensors deployed underwater. While previous ME-based underwater communication efforts have primarily focused on basic connectivity tests in controlled water-tank environments [29, 34, 35, 11], we validate BlueME in a natural lake and ocean, demonstrating its potential as a deployable communication system for practical marine robotics applications.
@@
We present a seamless integration of BlueME into pressure-compensated enclosures to ensure reliable underwater operations in real-time. We demonstrate that the enclosed system is compact and end-to-end portable for use on low-power embedded devices.
We validate the system through simulation as well as field experimental trials in freshwater (lake) and saltwater (ocean) environments, demonstrating its effectiveness for underwater communication. The system performs reliably when fully submerged and is unaffected by turbidity, line-of-sight obstacles, and shallow-water interference, which limit acoustic and optical systems.
Our field deployments reveal that we can achieve a high-fidelity communication link between mobile robots for up to meters with a power footprint of only - watt. These capabilities demonstrate its potential for robust and low-latency communication underwater.
@@ 2.1 Underwater Communication Technologies
Optical communication offers higher data rates and lower latency compared to acoustic methods [24, 25]. Notable systems include the BlueComm system, achieving 20 Mbps over distances up to 200 m [43]; the Ambalux system, providing 10 Mbps over 40 m [25]; and LUMA modems, supporting 10 Mbps across distances up to 50 m [44]. Underwater visible light communication (UVLC) has also been explored for extended ranges using wavelengths between 450–550 nm, which experience lower attenuation in water [26, 28]. However, optics are limited by line-of-sight requirements and are susceptible to scattering, absorption due to turbidity, and biofouling effects that degrade performance over time [45, 46].
Radio frequency (RF) electromagnetic communication underwater has historically been challenging due to high attenuation of EM waves in conductive media like seawater [14]. Attenuation increases with frequency, limiting effective communication to short distances [15]. However, RF communication offers advantages such as the ability to cross air-water and water-sediment boundaries, operation in turbid conditions, and non-line-of-sight capabilities [10, 47]. Recent research has revisited underwater RF communication using very low frequency (VLF, 3–30 kHz) and ultra-low frequency (ULF, 300–3000 Hz) bands to mitigate attenuation [29, 10]. Despite these advances, practical deployment is hindered by the need for large antennas and/or high transmission power, making it less suitable for mobile underwater applications [48, 12].
@@ 2.2 ME Antennas for Low-Power Communication
However, significant challenges remain in designing and fabricating compact, low-power ME antennas suitable for integration with mobile robotic platforms [54, 55]. Key issues include optimizing antenna performance, developing feasible deployment strategies, and characterizing communication properties in dynamic underwater conditions. This work aims to advance the field by developing a novel ME antenna array system tailored for marine robotics applications, addressing current gaps in practical deployment and performance evaluation.
@@ 3.1 ME Antenna Design: Theory and Simulation
To validate this theoretical analysis, we use COMSOL Multiphysics [61] to simulate a three-layer ME structure. As shown in Fig. 3, we evaluate the displacement magnitude of our antenna with a Metglas-PZT-Metglas thickness of -- m. By simulation, we predict the eigenfrequency for the bias-free fundamental displacement mode to be kHz.
Antenna Enclosure and Array Assembly. As shown in Fig. 4, each ME antenna is housed in a custom 3D-printed enclosure. The upper shell of the enclosure contains a pogo pin and two vertical AWG copper wires; the pogo pin is bonded to a designated antenna hole using a UV-curable cyanoacrylate adhesive. The AWG wires from the antenna are twisted and soldered to the AWG wires, which extend through the lower shell. The two shells are then enclosed and arranged in a grid within a frame that includes twenty N52 neodymium magnets. Each magnet is positioned from the antenna edge, all oriented with the same north–south polarity. The magnetic bias distribution across a representative antenna was measured as shown earlier in Fig. 3. Finally, the receiver and transmitter antennas are connected in series and parallel configurations, respectively.
3.2 Pressure Compensation & System Integration
@@ 4 System Integration & Experimental Setup
Our experimental analyses include (i) evaluation of the BlueME system’s magnetoelectric (ME) antenna communication characteristics, and (ii) performance assessment through robotics field trials. First, we validate the proposed antenna’s functionality using empirical data and simulation results, providing a basis for further investigation into optimal modulation techniques. We then integrate two BlueME systems on an autonomous surface vehicle (ASV) and an underwater remotely operated vehicle (ROV) to evaluate communication range and robustness, as illustrated in Fig. 6. To this end, we conduct a series of field experimental trials in open water to demonstrate the feasibility of the proposed BlueME system for robot-to-robot communication.
@@ 4.1 Setup: Measuring Antenna Impedance
(a) Freshwater experiments (lake trials)
(b) Saltwater experiments (ocean trials)
An impedance sweep from to was performed using a DAD3 board connected to an impedance measurement attachment and a BNC adapter. In these tests, the antennas were connected via short wires to a terminal and BNC oscilloscope probes. The impedance magnitude and phase were recorded for each antenna individually.
After individual tests, the antennas were wired together to form the transmitter and receiver arrays. The transmitter antennas were connected in parallel, while the receiver antennas were connected in series. The arrays were then tested as a whole, with impedance measurements taken under the same conditions as the individual antennas. The goal of this setup was to fabricate the ME antennas with sufficiently tight tolerances to validate that wiring the antennas in parallel or series effectively increases bandwidth and/or resonant response [62].
@@ 4.2 Setup: Underwater Communication Range
Lake Trials: We first conducted proof-of-concept experiments by evaluating the impedance of the proposed system in a water tank measuring m (L) m (W) m (D). The range tests were conducted using two ME antenna arrays in a freshwater lake. The receiver array was mounted on an underwater ROV (BlueROV2) and positioned at a depth of approximately below the surface, whereas the transmitter array was mounted on an ASV (BlueBoat) with the transmitter submerged approximately below the surface. The transmitter array was driven by a Class-D amplifier powered by a supply, which was generated using a boost converter connected to a LiFePO4 battery. The input signal to the amplifier varied from to . A small monitoring board was placed in line with the amplifier output, consisting of current sense resistors and an AC-coupled : voltage divider to measure the current through and voltage across the antenna array.
@@
For the range tests, the transmitter array conducted frequency sweeps from to at varying drive strengths. The receiver array collected data continuously, using the FFT peak hold feature of the WaveForms software to record the received signal amplitude across the frequency range. Initial tests were performed in freshwater at distances of , , , and . The ASV’s position during each trial was estimated via GPS with an accuracy of approximately and used to determine the transmitter–receiver separation (see Fig. 6). The maximum test distance was set to based on the available test area at the lake.
Ocean Trials: For more comprehensive and continuous testing, the experiments were repeated in saltwater off the west coast of Florida (Fig. 9). The Gulf test sites ranged from depths of approximately – (– m); range measurements were performed for horizontal distances. As shown in Fig. 6, the Tx (mounted on ASV) drifted continuously away, with the Tx assembly suspended 3 m below the water surface, while the Rx (mounted on ROV) operated near the ship at depths of -. GPS data from the ASV and the ship were used to determine the TX–RX separation, which was measured up to .
5 Performance Evaluation: Freshwater
@@ 5.1 Impedance Measurement Results
5.2 Freshwater Communication Performance
For the range tests, we initially anticipated that reactive near-field behavior would dominate the antenna response, resulting in rapid signal attenuation with distance. However, the measured results exhibited a slower decay rate than expected from purely reactive near-field coupling, with signals observed from to . These observations indicate that part of the transmitted energy propagated beyond the reactive near field, possibly through radiative or interface-guided modes along the water–air or water–seabed boundaries [63]. The maximum test distance in freshwater was set to based on the available test area; signal detection beyond this range was subsequently confirmed in ocean trials (see Sec. 6).
This performance can be attributed to several factors. First, the electromagnetic wavelength in water at is approximately , much shorter than in air due to water’s higher permittivity and conductivity. The wavelength in a conductive medium is given by , where is the phase constant. In conductive media, is influenced by permittivity, permeability, conductivity, and angular frequency. The reduced wavelength improves the radiation efficiency of electrically small antennas. Additionally, the ME antennas, when configured in arrays, are within each other’s near fields, effectively forming a larger equivalent magnetic dipole with increased radiation [11]. Antenna misalignment had a more pronounced effect at closer ranges. This can be attributed to near-field interactions causing destructive interference at close range and/or signal propagation along more diverse paths at greater distances.
Moreover, we found that the rate of attenuation decreased with increasing distance. This may be due to electromagnetic waves propagating along paths with lower attenuation (e.g., the lake bed, surface wave) [64] or due to reflections at shallow surfaces [15]. Unlike acoustic propagation, electromagnetic waves’ significantly higher propagation velocity in water results in wavelengths of orders of magnitude longer than acoustic waves at comparable frequencies. Consequently, typical path length differences produce smaller phase differences between multipath components, rendering interference effects negligible. Additionally, reflections from obstacles and natural features are minimal since objects of size comparable to or smaller than the electromagnetic wavelength tend to produce weak scattering, with incident waves primarily diffracting around or penetrating through such objects rather than producing strong reflections.
@@ 5.3 Analyzing the Theory of Operation
The performance characteristics of ME antenna arrays submerged underwater are not extensively explored in the literature. As such, we make several observations in our analysis.
Wavelength Changes. The antenna’s observed rate of signal drop-off appears to change across distances. This may be explained by the change in electromagnetic wavelength from air to fully submerged in water, which is also a more conductive medium. To this end, we calculate the expected wavelength of an electromagnetic plane wave in an underwater environment. As shown by [14], the electromagnetic propagation constant can be calculated as:
@@
Although the relative permittivity of water is significantly higher than air (:), the substantial decrease in wavelength (from in air to in water) results in a net increase in radiation resistance per antenna by a factor of approximately . This leads to an increase in the total radiated power per antenna by a similar factor.
@@
Increasing the number of antennas on the receiver side also increases the receiver sensitivity linearly, assuming all antennas are identical, i.e., a factor of improvement in the overall link budget. Consequently, using an array of antennas underwater compared to a single antenna in air theoretically yields an overall link budget improvement by a factor of , which corresponds to an increase of approximately dB – a factor not specifically discussed in the literature [11, 48, 29, 34, 35]. In practice, however, resonant frequency mismatches and increased mutual coupling at closer antenna spacings and higher drive strengths will degrade phase coherence within the transmitter array [62], reducing the actual achievable improvement from the array configuration.